if almonds sell at $1.20 per pound and walnuts sell for $.75 per pound, how many pounds of each must be used to make 45 pounds of a mixture to sell at $1.00 per pound?
Answer by checkley77(12844)

1.20x+.75(45-x)=1*45
1.20x+33.75-.75x=45
.45x=45-33.75
.45x=11.25
x=11.25/.45
x=25 pounds of almonds was used.
45-25=20 pounds of walnuts was used.
Proof:
1.20*25+.75*20=45
30+15=45
45=45